OpenStudy (anonymous):

HELP!! GIVING OUT MEDALS ;) NEED ANSWER NOW. Are the graphs of the lines in the pair parallel? Explain. y = 6x + 9 27x – 3y = –81 A.Yes, since the slopes are the same and the y-intercepts are different. B.Yes, since the slopes are the same and the y-intercepts are the same. C.No, since the slopes are different. D.No, since the y-intercepts are different.

OpenStudy (anonymous):

Alright so basically to figure out if two linear equations are parallel we put them in y=mx+b form and compare there slopes.

OpenStudy (anonymous):

The first equation is already in y=mx+b but we need to make the second one into the form

OpenStudy (anonymous):

To do this we must get y by itself

OpenStudy (anonymous):

First, subtract 27x to the other side to be \[-3y=-27x-81\]

OpenStudy (anonymous):

y is yet by itself and in order to do this we need to divide each side by -3 to get \[y=9x+27\]

OpenStudy (anonymous):

To see if these lines are parallel we look at the mx in y=mx+b If the mx are equal to each other they are parallel

OpenStudy (anonymous):

In this case... they are not.

OpenStudy (anonymous):

@gxldmouth Gucci?

OpenStudy (anonymous):

thaaank you .

OpenStudy (anonymous):

You're welcome ^-^

OpenStudy (anonymous):

wait is it C or D ?